If four times as many students pass a statistics course as those who fail and one statistics student is selected at random, what is the probability that the student will pass statistics?
Answer by sudhanshu_kmr(1152)


let number of failed students = x
number of passed students = 4x
total students = x + 4x = 5x
probability that a student will pass = 4x/5x = 4/5

answer : 4/5